12x^2+7x-12
Factor the Quadratic expression completely find the roots of the expression

Found 3 solutions by stanbon, solver91311, ewatrrr:
Answer by stanbon(75887)

12x^2+7x-12
Factor the Quadratic expression completely find the roots of the expression
Using the AC Method
12x+16-9x-12
4x(3x+4)-3(3x+4)
(4x-3)(3x+4) = 0
x = 3/4 or x = -4/3
======================
Cheers,
Stan H.

Answer by solver91311(24713)



You see, what I told you earlier wasn't so dumb. Using that information you should have been able to come up with the same solution that Stan showed you. I just wanted to give you the opportunity to use your head for something besides a hat rack rather than just handing you the solution so that you didn't have to think.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism



Answer by ewatrrr(24785)

 
Hi,
12x^2+7x-12
factoring
(4x -3 )(3x + 4) If x = 3/4 or x = -4/3 then (4x -3 )(3x + 4) = 0
F First terms 12x^2
O Outside terms +16x
I Inside terms -9x
L Last terms -12